# POUR: A Provably Optimal Method for Unlearning Representations via Neural Collapse

### v1

@@ Abstract
In computer vision, machine unlearning aims to remove the influence of specific visual concepts or training images without retraining from scratch. Studies show that existing approaches often modify the classifier while leaving internal representations intact, resulting in incomplete forgetting. In this work, we extend the notion of unlearning to the representation level, deriving a three-term interplay between forgetting efficacy, retention fidelity, and class separation. Building on Neural Collapse theory, we show that the orthogonal projection of a simplex Equiangular Tight Frame (ETF) remains an ETF in a lower dimensional space, yielding a provably optimal forgetting operator. We further introduce the Representation Unlearning Score (RUS) to quantify representation-level forgetting and retention fidelity. Building on this, we introduce POUR (Provably Optimal Unlearning of Representations), a geometric projection method with closed-form (POUR-P) and a feature-level unlearning variant under a distillation scheme (POUR-D). Experiments on CIFAR-10/100 and PathMNIST demonstrate that POUR achieves effective unlearning while preserving retained knowledge, outperforming state-of-the-art unlearning methods on both classification-level and representation-level metrics. Code will be released upon acceptance of the paper.
@@ 1 Introduction
The ability to selectively remove knowledge from trained models has become an increasingly important requirement for modern machine learning systems. Motivations include regulatory compliance with data protection laws (e.g., the “right to be forgotten”) [pipl2021, gdpr2016, ccpa2018], mitigating reliance on spurious correlations, and ensuring the safe deployment of large pre-trained models in sensitive domains such as autonomous driving and medical imaging. In these applications, models often need to forget outdated, biased, or privacy-sensitive visual data while retaining general visual understanding for reliable downstream use. Figure 1 illustrates this goal on the PathMNIST dataset, showing how our method erases the ”adipose” class while preserving the remaining categories.
@@
A growing literature on machine unlearning has explored how to make models forget a specific class, subset, or concept without retraining from scratch [bourtoule2021machine, golatkar2020eternal, chundawat2023zeroshot]. Previous work on machine unlearning has primarily focused on aligning the prediction probabilities on the forget and retain sets. This line of research, often referred to as weak unlearning [xu2023machineunlearningsurvey, golatkar2020forgetting], aims to ensure that the distributions of the final logits produced by the original and unlearned models are indistinguishable. However, recent studies [kim2025arewe] have questioned whether such methods truly forget the targeted information, as they often only perturb classifier logits while leaving the underlying feature representations largely unchanged. This shallow modification leaves residual information that can lead to privacy leakage [zhou2025decoupled]. This issue is particularly critical for deep vision encoders whose internal representations can still leak forgotten visual concepts through linear probing or feature inversion [kim2025arewe, golatkar2020eternal].
In parallel, theoretical advances have revealed that deep visual classifiers exhibit highly structured geometric behavior at convergence. The theory of Neural Collapse (NC) shows that class features concentrate around equidistant centroids and classifier weights align to form a simplex Equiangular Tight Frame (ETF) [papyan2020neuralcollapse]. This geometry provides a powerful lens for reasoning about class-level knowledge in image recognition: each class corresponds to a single ETF direction, and as we propose in this work, forgetting a class corresponds to removing its associated vector from the representation space. Previous work [kodge2024deep] has pursued heuristic realizations of ”projection as unlearning” through Singular Value Decomposition (SVD)-based decomposition in activation space. However, the method lacks geometric consistency and theoretical guaranties.
@@ 2.2 Practical Estimation of K
Due to the stochastic nature of training, comparing the feature distributions of the unlearned model and the retrained model is not straightforward. We require a representation measure that is robust to randomness, including random initialization, rotations of the feature basis, and uniform rescaling of feature magnitudes. Therefore, we adopt Centered Kernel Alignment (CKA) [kornblith2019similarity] as a practical estimator of representation similarity, for its invariance to scaling and rotation; additional justification is provided in Appendix.
@@ 5.2 Unlearning on CIFAR-10/100
Figure 4: t-SNE visualization of representation spaces after unlearning on CIFAR-10 and CIFAR-100. Each color denotes a retained class, with dark red points represent the forgotten class. The Gold panel shows the representation of the retrained model, serving as the ideal reference for successful unlearning. Structure of representations after POUR unlearning mostly resemble that of the retrained gold model.
@@ 0 More Related Work
Figure 1: Grad-CAM visualization on PathMNIST before and after unlearning. Each row shows a tissue class. Only after POUR unlearning, the Grad-CAM signal vanishes.

### v2

@@ Abstract
In computer vision, machine unlearning aims to remove the influence of specific visual concepts or training images without retraining from scratch. Studies show that existing approaches often modify the classifier while leaving internal representations intact, resulting in incomplete forgetting. In this work, we extend the notion of unlearning to the representation level, deriving a three-term interplay between forgetting efficacy, retention fidelity, and class separation. Building on Neural Collapse theory, we show that the orthogonal projection of a simplex Equiangular Tight Frame (ETF) remains an ETF in a lower dimensional space, yielding a provably optimal forgetting operator. We further introduce the Representation Unlearning Score (RUS) to quantify representation-level forgetting and retention fidelity. Building on this, we introduce POUR (Provably Optimal Unlearning of Representations), a geometric projection method with closed-form (POUR-P) and a feature-level unlearning variant under a distillation scheme (POUR-D). Experiments on CIFAR-10/100 and PathMNIST demonstrate that POUR achieves effective unlearning while preserving retained knowledge, outperforming state-of-the-art unlearning methods on both classification-level and representation-level metrics. Code is available at: https://github.com/ale256/representation_unlearning.
@@ 1 Introduction
The ability to selectively remove knowledge from trained models has become an increasingly important requirement for modern machine learning systems. Motivations include regulatory compliance with data protection laws (e.g., the “right to be forgotten”) [pipl2021, gdpr2016, ccpa2018], mitigating reliance on spurious correlations, and ensuring the safe deployment of large pre-trained models in sensitive domains such as autonomous driving and medical imaging. In these applications, models often need to forget outdated, biased, or privacy-sensitive visual data while retaining general visual understanding for reliable downstream use. Figure 1 illustrates this goal on the PathMNIST dataset, showing how our method erases the “adipose” class while preserving the remaining categories.
@@
A growing literature on machine unlearning has explored how to make models forget a specific class, subset, or concept without retraining from scratch [bourtoule2021machine, golatkar2020eternal, chundawat2023zeroshot, sepahvand2025selective]. Previous work on machine unlearning has primarily focused on aligning the prediction probabilities on the forget and retain sets. This line of research, often referred to as weak unlearning [xu2023machineunlearningsurvey, golatkar2020forgetting], aims to ensure that the distributions of the final logits produced by the original and unlearned models are indistinguishable. However, recent studies [kim2025arewe] have questioned whether such methods truly forget the targeted information, as they often only perturb classifier logits while leaving the underlying feature representations largely unchanged. This shallow modification leaves residual information that can lead to privacy leakage [zhou2025decoupled]. This issue is particularly critical for deep vision encoders whose internal representations can still leak forgotten visual concepts through linear probing or feature inversion [kim2025arewe, golatkar2020eternal].
In parallel, theoretical advances have revealed that deep visual classifiers exhibit highly structured geometric behavior at convergence. The theory of Neural Collapse (NC) shows that class features concentrate around equidistant centroids and classifier weights align to form a simplex Equiangular Tight Frame (ETF) [papyan2020neuralcollapse]. This geometry provides a powerful lens for reasoning about class-level knowledge in image recognition: each class corresponds to a single ETF direction, and as we propose in this work, forgetting a class corresponds to removing its associated vector from the representation space. Previous work [kodge2024deep] has pursued heuristic realizations of “projection as unlearning” through Singular Value Decomposition (SVD)-based decomposition in activation space. However, the method lacks geometric consistency and theoretical guaranties.
@@ 2.2 Practical Estimation of K
Due to the stochastic nature of training, comparing the feature distributions of the unlearned model and the retrained model is not straightforward. We require a representation measure that is robust to randomness, including random initialization, rotations of the feature basis, and uniform rescaling of feature magnitudes. Therefore, we adopt Centered Kernel Alignment (CKA) [kornblith2019similarity] as a practical estimator of representation similarity, for its invariance to scaling and rotation; additional justification is provided in Appendix 2.2.
@@ 5.2 Unlearning on CIFAR-10/100
Figure 4: t-SNE visualization of representation spaces after unlearning on CIFAR-10 and CIFAR-100. Each color denotes a retained class, with dark red points represent the forgotten class. The Gold panel shows the representation of the retrained model, serving as one possible reference for successful unlearning. Structure of representations after POUR unlearning mostly resemble that of the retrained gold model.
@@ 0 Related Work
1 Additional Baselines
We additionally report results (forget airplane class) for method [22], SalUn [fan2023salun], SCRUB [kurmanji2023towards] and SURE [sepahvand2025selective] in Tab. 6. POUR again achieves strongest performance across both classification- and representation-level metrics.
Table 6: More baselines on CIFAR-10.
[…]
Original 94.61 94.73 0.51 1.00 1.00 0.00
ProjUn [22] 86.49 20.40 0.76 0.34 0.89 0.76
SalUn 92.72 0.78 0.97 0.28 0.92 0.81
SCRUB 93.34 3.80 0.95 0.31 0.98 0.81
SURE 92.50 2.60 0.95 0.34 0.98 0.79
[…]